OpenStudy (anonymous):

Solve for the unknown variable. 3e^(7x)=72

OpenStudy (anonymous):

so far i got e^(7x)=24

OpenStudy (anonymous):

what is the next step?

OpenStudy (gorv):

now take log on both side

OpenStudy (gorv):

can u do that

OpenStudy (gorv):

after that use this property\[\ln a^b=b*\ln(a) \]

OpenStudy (anonymous):

\[x=\frac{ 1 }{ 7 } \ln (24)\] \[x=0.45400769005\]

OpenStudy (anonymous):

thanks you

OpenStudy (anonymous):

thanks @Jesstho.-.